# Parametric Autoresonance with Time-Delayed Control

### v1

@@ I Introduction
A specific type of nonlinear system in which the system becomes phase-locked with a driving force in an adiabatic manner, where the frequency of the forcing gradually changes over time. This phenomenon, widely known as autoresonance, leverages the system’s nonlinearity to maintain resonance and effectively increase the system’s amplitude over an extended period, despite changes in the system’s parameters sanjuanbook ; fajans2001autoresonant . Typically, when the parameters of a nonlinear system change, the resonant frequency also shifts, causing the system to become detuned from resonance. In autoresonance, however, the system remains naturally phase-locked with the driving force in an adiabatic manner, allowing for self-adjustment of both amplitude and frequency. Since the initial observation of autoresonance ms_livingstone , a plethora of studies have been conducted across various fields. The use and application of autoresonance have been explored in atomic physics meerson1990strong ; liu1995nonlinear , plasma physics fajans1999autoresonant1 ; fajans1999autoresonant2 ; andresen2011autoresonant ; baker2015electron , nonlinear wave interactions friedland1998autoresonance3 ; friedland1998autoresonant4 ; yaakobi2013complete , planetary dynamics malhotra1999migrating ; friedland2001migration ; lanza2022tidal , and fluid dynamics friedland1999control . The theoretical framework of autoresonance has been extensively developed over the past few years. Some of the foundational work has been carried out by L. Friedland et al. friedland1997variational ; fajans2001dampingeffect ; fajans1999collective ; fajans2000secondharmonic ; nakar1999passage ; barth2014quantum . Studies can be found on asymptotic analysis and stability of autoresonance by L.A. Kalyakin et al. kalyakin2008asymptotic ; kalyakin2013stability . Kovaleva et al. kovaleva2013limiting have contributed an array of excellent works, including the limiting phase trajectory description of autoresonance and investigations of autoresonance in nonlinear coupled chains manevitch2016autoresonant ; kovaleva2018autoresonance ; kovaleva2016autoresonance . Additionally, R. Chacón et al. chacon2005energy ; chacon2010universal ; chacon2008breakdown have developed an energy-based theory and conducted notable research on chaos and the breakdown of autoresonance . While autoresonance induced by external forcing has garnered significant attention, relatively few studies have focused on parametric autoresonance khain2001parametric ; assaf2005parametric ; kiselev2007capture ; friedland2016parametric .
@@
In this article, we have investigated the effect of time delay on a parametric autoresonant system. The control of resonance response whether deterministic or stochastic, in various nonlinear systems with time delay has proven to be an efficient strategy over the years hu1998resonances ; mei2009effects ; jeevarathinam2011theory ; cantisan2020delay ; zakharova2017time ; maccari2003vibration . In a recent work, the effect of delay has been discussed in a externally driven autoresonant array of Duffing-Ueda oscillator chacon2024 . Although time delay has been employed in systems with externally applied chirped forcing, its role in parametric autoresonant systems warrants attention due to its potential for effective applications across diverse fields. Driven by an interest in parametric autoresonance, we aim to investigate the effect of time delay in such systems, an aspect that has received limited attention in the existing literature, to the best of our knowledge. Our findings demonstrate that the delay strength can serve as a powerful mechanism to control the growth of the autoresonant system.
@@ II The model
where the parametric forcing $h\cos\nu$, where the frequency now depends on time and is defined as $\dot{\nu}=2\omega-2\mu t$, with $\mu$ representing the chirp rate. Here $\gamma$ is the damping constant, and $\omega_{0},\alpha$ are the coefficients of linear and nonlinear stiffness, respectively. The model incorporates a constant delay $\bar{\beta}$ with strength $k$. Now introducing a dimensionless time $\omega t=\tau$, we can rewrite Eq. (1) as
@@ II.1 Slow-flow and amplitude equation
$\sigma=\frac{1}{\omega^{2}}\bigg{[}(2\mu t+k\cos(\omega\bar{\beta})-\frac{3\alpha\hat{A}^{2}}{4})\pm\sqrt{\omega^{4}h^{2}-4(\omega\gamma-k\sin(\omega\bar{\beta}))^{2}}\bigg{]}.$ (12)
@@
$\hat{A}=\sqrt{\frac{4}{3\alpha}\bigg{[}(2\mu t+k\cos(\omega\bar{\beta})\pm\sqrt{\omega^{4}h^{2}-4(\omega\gamma-k\sin(\omega\bar{\beta}))^{2}}-\frac{\omega^{2}-\omega_{0}^{2}}{\epsilon}\bigg{]}}.$ (13)
Equation (13) shows the amplitude growth over time, as shown in Fig. 1. In the following section, we analyze the stability of the system, showing that for stable, autoresonant amplitude growth, the delay strength must exceed a specific threshold. Below this threshold, the fixed points become unstable, making analytical expressions infeasible; however, numerical analysis can still be performed.
@@ II.2 Stability and the critical delay strength
For the parameter values $\gamma=0.002$, $\omega=\omega_{0}=0.5$,$\sigma=0.1$,$h=0.0002$ and $\alpha=0.08$,we have plotted the analytical expression in Eq. (13). For $\bar{\beta}=0.1$, the threshold value is $k_{th}=0.02$ and for $\bar{\beta}=0.3$, the threshold is $k_{th}=0.0066$ according to Eq. (18). When these thresholds are exceeded, autoresonant growth is observed. In the subsequent section, we provide numerical demonstrations showing that autoresonant behavior is triggered when this critical threshold is crossed and diminishes below it. In the following section, we numerically demonstrate how autoresonant behavior is triggered above this critical threshold and diminishes below it.
@@ III Numerical results
After conducting an analytical study of the system, we believe it is crucial to validate these results through numerical testing. To achieve this, we have calculated several key aspects: the oscillations along the $x$-axis, the corresponding phase-space orbits, the maximum amplitude of these oscillations, which provides the envelope of the oscillations, and the evolution of this maximum over time, derived using an exponential fit of the form $x=ae^{bt}$. In our simulations, we have fixed $\gamma=0.002,\mu=0.00001,\omega=0.5,\alpha=0.03$ and we have varied $h,k,$ and $\bar{\beta}$.
Contrarily to the theoretical analysis, shown in Fig. 1(a), for the case where $k\leq 0.02$, the numerical results indicate a negative trend in the maximum amplitude of the oscillations. In fact, numerically the threshold value of $k_{th}=0.0201$, even if it is a little bit higher, it is still in good agreement with the theoretical analysis. In Fig. 2(a), Fig. 2(c), and Fig. 2(e) this is evident both visually in the decay of the oscillation envelope and mathematically in the negative sign of the exponential parameter $b$. Conversely, when $k=0.024$, the behavior changes; in Fig. 2(b), Fig. 2(d), and Fig. 2(f) the oscillations exhibit a growing trend, confirmed graphically and by a positive $b$, which signifies that autoresonance has been initiated for this parameter value. These numerical findings are in strong agreement with the predictions made by our analytical model.

### v2

@@ I Introduction
We consider a specific type of nonlinear system in which the system becomes phase-locked with a driving force in an adiabatic manner, where the frequency of the forcing gradually changes over time. This phenomenon, widely known as autoresonance, leverages the system’s nonlinearity to maintain resonance and effectively increase the system’s amplitude over an extended period, despite changes in the system’s parameters sanjuanbook ; fajans2001autoresonant . Typically, when the parameters of a nonlinear system change, the resonant frequency also shifts, causing the system to become detuned from resonance. In autoresonance, however, the system remains naturally phase-locked with the driving force in an adiabatic manner, allowing for self-adjustment of both amplitude and frequency. Since the initial observation of autoresonance ms_livingstone , a plethora of studies have been conducted across various fields.
The theoretical framework of autoresonance has been extensively developed over the past few years. Some of the foundational work has been carried out by L. Friedland et al. friedland1997variational ; fajans2001dampingeffect ; fajans1999collective ; fajans2000secondharmonic ; nakar1999passage ; barth2014quantum . Studies can be found on asymptotic analysis and stability of autoresonance by L.A. Kalyakin et al. kalyakin2008asymptotic ; kalyakin2013stability . Kovaleva et al. kovaleva2013limiting have contributed an array of works, including the limiting phase trajectory description of autoresonance and investigations of autoresonance in nonlinear coupled chains manevitch2016autoresonant ; kovaleva2018autoresonance ; kovaleva2016autoresonance . Additionally, R. Chacón et al. chacon2005energy ; chacon2010universal ; chacon2008breakdown have developed an energy-based theory and conducted a research on chaos and the breakdown of autoresonance . While autoresonance induced by external forcing has garnered significant attention, relatively few studies have focused on parametric autoresonance khain2001parametric ; assaf2005parametric ; kiselev2007capture ; friedland2016parametric .
@@
Here, we have investigated the effect of time delay on a parametric autoresonant system. The control of resonance response whether deterministic or stochastic, in various nonlinear systems with time delay has proven to be an efficient strategy over the years hu1998resonances ; mei2009effects ; jeevarathinam2011theory ; cantisan2020delay ; zakharova2017time ; maccari2003vibration . In a recent work, the effect of delay has been discussed in a externally driven autoresonant array of Duffing-Ueda oscillator chacon2024 . Although time delay has been employed in systems with externally applied chirped forcing, its role in parametric autoresonant systems warrants attention due to its potential for effective applications across diverse fields. Chirped forcing refers to time-dependent parametric excitation where the driving frequency varies gradually over time. This gradual change allows phase-locking between the system and the driving force, a mechanism central to sustaining autoresonance despite parameter shifts. While autoresonance driven by external forces has been extensively studied, understanding the dynamics of parametric forcing with time-delayed feedback offers new opportunities for controlling resonance stability and amplitude growth.
In this work, we aim to analyze how time-delay feedback influences the stability of parametric autoresonance. Then, to derive a critical threshold for delay strength that governs the transition from decaying to growing oscillations. finally we validate these analytical predictions using numerical simulations. The ability to tune resonance behavior through time delay broadens potential control strategies in mechanical, electrical, and fluid systems. Driven by an interest in parametric autoresonance, we aim to investigate the effect of time delay in such systems, an aspect that has received limited attention in the existing literature, to the best of our knowledge. Our findings demonstrate that the delay strength can serve as a mechanism to control the growth of the autoresonant system.
@@ II The model
where the parametric forcing $h\cos\nu$, where the frequency now depends on time and is defined as $\dot{\nu}=2\omega-2\mu t$.The term $\nu(t)$ in the model evolves as specified before with $\mu$ representing the chirp rate, enabling phase-locking and continuous resonance. This follows from the definition of chirped forcing that refers to time-dependent parametric excitation where the driving frequency varies gradually. Here $\gamma$ is the damping constant, and $\omega_{0},\alpha$ are the coefficients of linear and nonlinear stiffness, respectively. The model incorporates a constant delay $\bar{\beta}$ with strength $k$. Now introducing a dimensionless time $\omega t=\tau$, we can rewrite Eq. (1) as
@@
The equation of the model (Eq. 1) models a parametric autoresonant system with delayed feedback, a generalization of Duffing-Ueda-like systems to incorporate time-delay effects. Time delay is critical for practical implementations where control and feedback loops introduce lag due to sensing and actuation dynamics. Real-world analogs include mechanical systems like pendulums with modulated pivots or beams with variable stiffness, as well as LC circuits where capacitance or inductance varies over time. In these contexts, sensors and controllers facilitate parametric forcing and delay management.
@@ II.1 Slow-flow and amplitude equation
$\sigma=\frac{1}{\omega^{2}}\bigg{[}\bigg{(}2\mu t+k\cos(\omega\bar{\beta})-\frac{3\alpha\hat{A}^{2}}{4}\bigg{)}\pm\bigg{(}\omega^{4}h^{2}-4(\omega\gamma-k\sin(\omega\bar{\beta}))^{2}\bigg{)}^{1/2}\bigg{]}.$ (12)
@@
$\hat{A}=\bigg{[}\frac{4}{3\alpha}\bigg{(}(2\mu t+k\cos(\omega\bar{\beta})\pm\sqrt{\omega^{4}h^{2}-4(\omega\gamma-k\sin(\omega\bar{\beta}))^{2}}-\frac{\omega^{2}-\omega_{0}^{2}}{\epsilon}\bigg{)}\bigg{]}^{1/2}.$ (13)
Equation (13) shows the amplitude growth over time, as shown for the parameter values $\gamma=0.002$, $\omega=\omega_{0}=0.5$, $\sigma=0.1$, $h=0.0002$ and $\alpha=0.08$, in Fig. 1. In the following section, we analyze the stability of the system, showing that for stable, autoresonant amplitude growth, the delay strength must exceed a specific threshold. Below this threshold, the fixed points become unstable, making analytical expressions infeasible; however, numerical analysis can still be performed.
@@ II.2 Stability and the critical delay strength
For $\bar{\beta}=0.1$, $\gamma=0.002$, and $\omega=0.5$ the threshold value is $k_{th}=0.02$ and for $\bar{\beta}=0.3$, the threshold is $k_{th}=0.0066$ according to Eq. (18). When these thresholds are exceeded, autoresonant growth is observed. In the subsequent section, we provide numerical demonstrations showing that autoresonant behavior is triggered when this critical threshold is crossed and diminishes below it. In the following section, we numerically demonstrate how autoresonant behavior is triggered above this critical threshold and diminishes below it.
@@ III Numerical results
We validate the analytical predictions by conducting numerical simulations across a range of system parameters. Specifically, we examine the amplitude evolution and phase-space dynamics under varying delay strength $k$ and time-delay $\bar{\beta}$ values. We quantify the autoresonance onset by fitting the oscillation envelope using an exponential model $x=ae^{bt}$. We used an exponential fit to analyze the behavior of the oscillation amplitude and its envelope. The parameter $b$ in the fit indicates different growth regimes: $b>0$ signifies increasing amplitude, $b<0$ represents decreasing amplitude, and $b\approx 0$ denotes constant amplitude growth. This approach effectively distinguishes between growing, decaying, and steady-state oscillations. In our simulations, we have fixed $\gamma=0.002,\mu=0.00001,\omega=0.5,\alpha=0.03$ and we have varied $h,k,$ and $\bar{\beta}$. An additional study examining the impact of varying the initial conditions on the phenomenon was conducted, with no qualitative differences in the outcomes observed.
Contrarily to the theoretical analysis, shown in Fig. 1(a), for the case where $k\leq 0.02$, the numerical results indicate a negative trend in the maximum amplitude of the oscillations. In fact, numerically the threshold value of $k_{th}=0.0201$, even if it is a little bit higher, it is still in good agreement with the theoretical analysis. In Fig. 2(a), Fig. 2(c), and Fig. 2(e) this is evident both visually in the decay of the oscillation envelope and mathematically in the negative sign of the exponential parameter $b$. Conversely, when $k=0.024$, the behavior changes; in Fig. 2(b), Fig. 2(d), and Fig. 2(f) the oscillations exhibit a growing trend, confirmed graphically and by a positive $b$, which signifies that autoresonance has been initiated for this parameter value. Worth to be mentioned is that a similar study has been conducted for the case of Fig. 1(b). Also, for $\bar{\beta}=0.3$ the numerical threshold value of the parameter $k$ is similar to the value found from Eq. 18, being the numerical value $k_{th}\approx 0.0068$. These numerical findings are in strong agreement with the predictions of the threshold value of $k$ made by our analytical results, obtained from Eq. 18.
@@ IV Conclusions
In summary, we have demonstrated that the presence of time delay significantly influences the stability and growth of a parametric autoresonant system. Using both analytical techniques and numerical simulations, we have identified a critical delay strength $k_{\text{th}}$ that marks the onset of sustained amplitude growth. Our numerical simulations confirm that for values of $k>k_{\text{th}}$, the oscillation amplitude increases, while for $k<k_{\text{th}}$, it decays, aligning well with the theoretical predictions derived from Eq. (18). Additionally, we found that the interplay between $k$ and the delay time $\bar{\beta}$ exhibits a more complex relationship than initially predicted, highlighting areas for future exploration in the control of parametric resonance.
These findings pave the way for practical applications where precise resonance control is needed, such as mechanical systems with variable stiffness or feedback-controlled electronic circuits, to further broaden the understanding of delay-induced resonance phenomena.